# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Coding

### v1

@@ Abstract.
Generative AI (GenAI) tools have already started to transform software development practices. Despite their utility in tasks such as writing code, the use of these tools raises important legal questions and potential risks, particularly those associated with copyright law. In the midst of this uncertainty, this paper presents a study jointly conducted by software engineering and legal researchers that surveyed 574 GitHub developers who use GenAI tools for development activities. The survey and follow-up interviews probed the developers’ opinions on emerging legal issues as well as their perception of copyrightability, ownership of generated code, and related considerations. We also investigate potential developer misconceptions, the impact of GenAI on developers’ work, and developers’ awareness of licensing/copyright risks. Qualitative and quantitative analysis showed that developers’ opinions on copyright issues vary broadly and that many developers are aware of the nuances these legal questions involve. We provide: (1) a survey of 574 developers on the licensing and copyright aspects of GenAI for coding, (2) a snapshot of practitioners’ views at a time when GenAI and perceptions of it are rapidly evolving, and (3) an analysis of developers’ views, yielding insights and recommendations that can inform future regulatory decisions in this evolving field.
@@ 1. Introduction
Generative AI (GenAI) tools have become widely adopted across many different domains, including software engineering (SE)  (R. Li et al., 2023; Ebert and Louridas, 2023; Rajbhoj et al., 2024; Krill, [n. d.]). Several GenAI coding assistants, including GitHub’s Copilot (GitHub, [n. d.]), Tabnine (Tabnine, [n. d.]), Codeium (codeium, [n. d.]), and Cody (cody, [n. d.]), as well as general purpose tools such as ChatGPT (OpenAI, [n. d.]), Claude (Anthropic Models, [n. d.]), and Gemini (Gemini Team et al., 2023), have become readily accessible, either as IDE extensions or standalone applications, enabling developers to perform many coding tasks with little effort, including automated code completion, summarization, and debugging. A 2024 Stack Overflow survey reported that 76 percent of 65,000 participating developers were using or planning on using AI coding tools (May, [n. d.]); more than one million developers used Copilot (GitHub, [n. d.]) in its first year (Yao, 2023). The popularity of these tools stems from the fact that they are easy to access and can generate diverse content with relatively little effort, thus helping users accomplish many tasks more efficiently (Liang et al., 2024; Barke et al., 2023).
@@
Alongside these benefits, however, using GenAI tools can introduce various issues, including bias or discrimination (Ferrara, 2023), security threats (Majdinasab et al., 2024a), and compromise of private information (Gupta et al., 2023). GenAI can also give rise to legal risk related to intellectual property concerns, including copyright infringement (Generative Artificial Intelligence and Copyright Law, [n. d.]). Many legal questions regarding using GenAI tools remain unanswered, and any answers will likely vary among jurisdictions and across cases.
@@
As of this writing, there are several pending lawsuits against high-profile providers of GenAI tools including OpenAI (Authors Guild v. OpenAI Inc., 2023), StabilityAI (get, 2023), and Midjourney (Andersen v. Stability AI Ltd., [n. d.]), as well as active consideration by the U.S. Copyright Office and other governmental entities around the world of the copyright issues raised by the use of GenAI (Office, 2024, 2023; Council of the European Union, 2024b). These matters implicate several considerations, including the extent to which works associated with the use of GenAI, including prompts and models, are protected by copyright and, if so, who owns the rights to such works; whether the use of open source software and other protected works as training data results in license violations or otherwise constitutes copyright infringement; and whether output resulting from the use of GenAI that is similar to preexisting work constitutes copyright infringement.
Because GenAI tools are widely used, many software developers and organizations are likely engaging in activities that implicate potential legal risks despite not having the legal training or legal advice needed to evaluate these risks. Regulation of this space– whether by governmental entities or corporate entities– should, therefore, not proceed without an awareness of the preexisting views of stakeholder groups with a vested interest in these issues. In this paper, we aim to further the discourse surrounding GenAI in the SE space and inform future decisions by identifying developers’ views on (1) using GenAI tools for software development tasks and (2) associated legal concerns, focusing specifically on copyright law.
@@
This paper presents a study conducted by a joint team of SE and legal researchers that surveyed 574 software developers worldwide who use GenAI tools for coding tasks (particularly code generation). Through an online survey and follow-up interviews, we probed the developers’ opinions on potential emerging legal issues, the perception of what is copyrightable, ownership of generated code, and related considerations. We also sought to understand potential developer misconceptions, assess the impact of GenAI on their work, and evaluate their awareness of licensing/copyright risks. Using qualitative and quantitative methods to analyze survey responses, we found that developer opinions on copyright issues vary broadly, particularly on the topic of model output ownership, and that many developers are aware of the nuances and complications associated with answering these complex legal questions. We discuss the results of our study, interpret the findings under the background of U.S. law, and offer insights for future work.
In summary, we make the following contributions: (1) a comprehensive survey of 574 developers worldwide on the licensing and copyright aspects of GenAI for coding, (2) a snapshot of practitioners’ views at a time when GenAI and perceptions of it are rapidly evolving, and (3) a rigorous analysis of developers’ views on these topics, resulting in insights and recommendations that can help inform future policy decisions in this evolving field. We make available our survey, results, and other artifacts for transparency and validation in an online replication package (ano, 2024).
2. Background
@@
Unless it is in the public domain, any textual work is generally protected by copyright in the U.S. so long as it constitutes an original work of authorship and is fixed in a tangible medium of expression (17 U.S.C. § 102, a). Similar provisions exist in the intellectual property legislation of other countries, including those in the European Union (Commission, 2024). The owner of a copyright—which may be an individual author, joint authors, or a corporate author as the owner of a work for hire—has several exclusive rights under U.S. copyright law, including the rights to reproduce the work in copies, to create derivative works, and to distribute the work, which also includes the right to authorize others to engage in such activities (17 U.S.C. § 106, [n. d.]).
The use of GenAI implicates these rights in several ways that are currently under consideration by the U.S. Copyright Office and the courts. The U.S. Copyright Office opined in February 2023—reviewing an application for copyright registration—that images generated by Midjourney were not the product of human authorship and were therefore ineligible for copyright registration (cop, 2023). The U.S. Copyright Office maintains this position (Office, 2023). There are also multiple pending cases alleging that the inclusion of copyrighted works in the training data of large language models (LLMs) and the generation of material that is substantially similar to those works constitutes infringement. For example, New York Times Co. v. Microsoft Corporation (The New York Times Company v. Microsoft Corporation, [n. d.]), filed in December 2023, alleges that ChatGPT uses the newspaper’s content as training data, memorizes that data and then not only provides outputs that are nearly identical to that content but hallucinates content that is then incorrectly attributed to the Times. Another example is Doe v. GitHub (DOE 1 et al v. GitHub, [n. d.]), filed in November 2022, which alleges that Copilot violates the copyright licenses of the plaintiffs’ open source software by using that code as training data and generating code that is a near-identical reproduction of the plaintiffs’ code but without adhering to the terms of the associated licenses. Yet another lawsuit, filed by a group of authors, was brought against the companies Meta, Microsoft, and Bloomberg, alleging that the authors’ works were used for training without their permission (Brittain, 2023).
@@
The resolution of these and similar cases depends on both technological and legal questions. On the technological side, liability may depend on whether authors can successfully show that their works are indeed included in a particular LLM’s training data, as well as whether they can show that those works are copied for any length of time during the training process (or whether the training simply involves developing model parameters from the training data without reproducing it). On the legal side, liability may depend on whether and how the generated content is similar to any particular author’s work. For example, generated content that is similar to previous works only with respect to the underlying concept or function is not infringing because the protections granted by U.S. copyright law for an original work of authorship do not extend to any “idea, procedure, process, system, method of operation, concept, principle, or discovery, regardless of the form in which it is described, explained, illustrated, or embodied in such work” (17 U.S.C. §102, b).
Additionally, some legal scholars argue that the use of a work protected by copyright as part of a model’s training data may be deemed a fair use under U.S. law (Lemley and Casey, 2020), similar to how the use by Google of copyrighted material as the database for its Google Books search engine was held to be fair use by the U.S. Court of Appeals for the Second Circuit in 2015 (Authors Guild, Inc. v. Google, Inc., 2015). Part of this consideration under U.S. law is likely to be whether the challenged use has a “transformative” purpose or is considered to be an unlawful derivative work; the 2021 U.S. Supreme Court case of Google LLC v. Oracle America, Inc. (Google LLC v. Oracle America, Inc., [n. d.]) and the 2023 U.S. Supreme Court case of Andy Warhol Foundation for the Visual Arts, Inc. v. Goldsmith (war, 2023) offer contrasting examples.
@@
Legislation both within and outside the U.S. also has implications for GenAI and copyright. Various U.S. states have attempted to enact legislation to govern GenAI (colorado-bill, 2024; california-regulation, 2024). In March 2024, the European Parliament approved the EU AI Act (Council of the European Union, 2024b), which, among other things, requires the providers of general-purpose AI models to “put in place a policy to comply with Union law on copyright and related rights, and in particular to identify and comply with, including through state-of-the-art technologies, a reservation of rights expressed pursuant to Article 4(3) of Directive (EU) 2019/790” (Council of the European Union, 2024a). In 2023, an early announcement of the EU Global Principles for Artificial Intelligence that led to the EU AI Act encouraged some organizations, including French media organizations France Médias Monde and TF1, to forbid companies such as OpenAI from mining their data (Karathanasis, 2023); that year also saw a temporary ban on ChatGPT in Italy (McCallum, 2023). Even countries like Japan, which initially took more relaxed stances, affirming that training AI models on protected content was not a violation of copyright law (Lanz, [n. d.]), have started to consider additional restrictions designed to protect copyright holders (Shinbun, 2023). Gervais et al. (Gervais et al., 2024) provide an in-depth overview of the current copyright landscape around the globe.
@@
Additional work relevant to GenAI and copyright issues focuses on research that may help answer some of the above technological questions. For example, techniques exist to potentially reveal whether a machine learning model was trained on a certain data record, such as membership inference attacks (Duan et al., 2024; Hu et al., 2022; Majdinasab et al., 2024b; Zhang and Li, 2023; Z. Yang et al., 2023). Such attacks can extract training data from the target model’s output, including private and personal information (Carlini et al., 2021). Memorization of training data is a known issue, including in models trained on code (Yang et al., 2024), with research suggesting that typical measures aimed at preventing verbatim extraction of training data are not sufficient to prevent the data from being obtained with other methods, such as by utilizing style-transfer prompts (Ippolito et al., 2023). Tools like CopyrightCatcher (CopyrightCatcher, [n. d.]) have been developed to detect the presence of copyrighted material in model output, evidencing that this is an issue of concern to developers.
Our work builds on this related work by contributing an in-depth exploration of developers’ perceptions of the copyright issues surrounding the use of AI-based code-generation tools, both in the U.S. and elsewhere around the world. Understanding these perspectives allows future regulation in this area to be informed by beliefs on the ground rather than in a vacuum.
@@ 3. Study Methodology
The goal of this study is to investigate practitioners’ perceptions about licensing and copyright issues related to the use of GenAI technology for software development. (Although we use the term “software development” throughout the paper, our study primarily focuses on the use of GenAI for code generation as the main task that can lead to copyright and licensing issues.) The primary quality focus is the mitigation of legal consequences due to copyright and licensing violations that arise when developing software with the assistance of GenAI. The context consists of 574 developers of open-source projects hosted on GitHub, whom we surveyed via an online survey and follow-up interviews.
@@
    What other legal concerns do developers anticipate as the use of GenAI increases? There are a plethora of other potential legal issues to consider as GenAI becomes more widely adopted, including data privacy, generation of malicious content, and tort liability. In this RQ, we analyze developers’ perspectives on these topics.
@@ 3.1. Survey design
To design our survey, we followed general guidelines (Groves et al., 2009) as well as SE-specific best practices (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). The final questionnaire went through multiple iterations of review and improvements from six SE researchers and one law researcher. We carefully formulated the questions and ensured they were written clearly and concisely to avoid biasing respondents. We also conducted a pilot study with graduate students from our research lab who actively use GenAI tools. Based on their feedback, we further refined our questionnaire, improving questions that were ambiguous or hard to understand.
@@
  • The section on “Current use of GenAI for coding” (12 questions) asked about respondents’ experience using GenAI tools for code generation, including the tools they used; the benefits and challenges of using such tools; the development tasks for which they used the tools; organizational policies of using such tools; and procedures to document tool usage, among other aspects.
  • The section on “Understanding and Perception of Copyright Issues” (14 questions) asked about developers’ perceptions of copyright issues surrounding GenAI, including the use of their code in model training without permission; having models produce code similar to theirs; reusing without permission a prompt that they created; ownership and copyrightability of generated code; whether the use of source code in training models should require attribution and/or monetary compensation; and the effects of these concerns on developers’ workflows.
  • The “Demographics” section (8 questions) asked about respondents’ experience with SE and AI, whether they develop open source or proprietary software, their location, and their primary programming languages, among other questions.
The survey was designed to be completed in about 15 to 20 minutes and consisted mostly of multiple-choice questions (26), complemented by open-ended questions (7) that allowed participants to elaborate on their responses. Not every participant saw every question, as we included logic in our survey to ask only those questions relevant to a participant’s indicated experience. The full survey text and a brief description are found in our replication package (ano, 2024).

### v3

# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Software Development

@@ Abstract.
Despite the utility that Generative AI (GenAI) tools provide for tasks such as writing code, the use of these tools raises important legal questions and potential risks, particularly those associated with copyright law. As lawmakers and regulators engage with those questions, the views of users can provide relevant perspectives. In this paper, we provide: (1) a survey of 574 developers on the licensing and copyright aspects of GenAI for coding, as well as follow-up interviews; (2) a snapshot of developers’ views at a time when GenAI and perceptions of it are rapidly evolving; and (3) an analysis of developers’ views, yielding insights and recommendations that can inform future regulatory decisions in this evolving field. Our results show the benefits developers derive from GenAI, how they view the use of AI-generated code as similar to using other existing code, the varied opinions they have on who should own or be compensated for such code, that they are concerned about data leakage via GenAI, and much more, providing organizations and policymakers with valuable insights into how the technology is being used and what concerns stakeholders would like to see addressed.
@@ 1. Introduction
Generative AI (GenAI) tools have become widely adopted across many different domains, including software engineering (SE)  (R. Li et al., 2023; Ebert and Louridas, 2023; Rajbhoj et al., 2024; Krill, [n. d.]). Several GenAI coding assistants, including GitHub’s Copilot (GitHub, [n. d.]), Tabnine (Tabnine, [n. d.]), Codeium (codeium, [n. d.]), and Cody (cody, [n. d.]), as well as general purpose tools such as ChatGPT (OpenAI, [n. d.]), Claude (Anthropic Models, [n. d.]), and Gemini (Gemini Team et al., 2023), have become readily accessible, either as IDE extensions or standalone applications, enabling developers to perform many coding tasks with little effort, including automated code completion, summarization, and debugging. A 2024 Stack Overflow survey reported that 76 percent of 65,000 participating developers were using or planning on using AI coding tools (May, [n. d.]); more than one million developers used Copilot (GitHub, [n. d.]) in its first year (Yao, 2023). The popularity of these tools stems from the fact that they are easy to access and can generate diverse content with relatively little effort, thus helping users perform many tasks more efficiently (Liang et al., 2024; Barke et al., 2023). Specifically, these emerging GenAI tools can help developers with various software engineering tasks, including code autocompletion, summarization, and debugging.
@@
However, in addition to these benefits, the use of GenAI tools can introduce various issues, including bias or discrimination (Ferrara, 2023), security threats (Majdinasab et al., 2024a), and compromise of private information (Gupta et al., 2023). GenAI can also pose legal risks related to intellectual property concerns, including copyright infringement (Generative Artificial Intelligence and Copyright Law, [n. d.]). Many legal questions regarding the use of GenAI tools remain unanswered, and any answers will likely vary among jurisdictions and across cases.
@@
As of this writing, there are several pending lawsuits against high-profile providers of GenAI tools including OpenAI (Authors Guild v. OpenAI Inc., 2023), StabilityAI (get, 2023), and Midjourney (Andersen v. Stability AI Ltd., [n. d.]), as well as active consideration by the U.S. Copyright Office and other governmental entities around the world of the copyright issues raised by the use of GenAI (Office, 2024a, 2023; Council of the European Union, 2024b; Office, 2025). These matters implicate several considerations, including the extent to which works associated with the use of GenAI, including prompts and models, are protected by copyright and, if so, who owns the rights to such works; whether the use of open source software and other protected works as training data results in license violations or otherwise constitutes copyright infringement; and whether output resulting from the use of GenAI that is similar to preexisting work constitutes copyright infringement.
Because any regulation of this space– whether by governmental entities or corporate entities– should take account of current and future practices, we aim in this paper to identify developers’ views on (1) using GenAI tools for software development tasks and (2) associated legal concerns, focusing specifically on copyright law. This is motivated by recent solicitations from the United States Copyright Office requesting the opinions of experts, organizations, and individuals (e.g., developers) for active consideration in the development of future regulation (Office, 23).
@@
This paper presents a study conducted by a joint team of SE and legal researchers that surveyed 574 software developers worldwide who use GenAI tools for software development tasks (particularly code generation). Through an online survey and follow-up interviews, we investigated developers’ opinions on potential emerging legal issues, the perception of what is copyrightable, ownership of generated code, and related considerations. We also sought to understand potential developer misconceptions, assess the impact of GenAI on their work, and assess their awareness of licensing/copyright risks.
Using qualitative and quantitative methods to analyze survey responses, we found that developer opinions on copyright issues vary broadly, particularly on the topic of model output ownership, and that many developers are aware of the nuances and complications associated with answering these complex legal questions. We discuss the results of our study, interpret the findings under the background of U.S. law, and offer insights for future work. We focus on developers rather than other stakeholders, such as compliance officers and policymakers, because developers are often the ones who directly interact with GenAI during software development. We survey an international population of developers because software development is a global activity, particularly for open-source projects. Additionally, large technology companies regularly employ individuals from various countries (and jurisdictions) around the world. Given this, it is appropriate to consider the thoughts, opinions, and perspectives of developers regardless of their geographic location.
We make available our survey, results, and other artifacts for transparency and validation in an online replication package (ano, 2024).
2. Background and Related Work
This intersectional field of copyright and generative AI relies on concepts drawn both from the legal domain and from software engineering. We discuss these concepts in this section as well as prior work related to our study.
2.1. Legal background
@@
Unless it is in the public domain, computer code is generally protected by copyright in the U.S. so long as it constitutes an original work of authorship and is fixed in a tangible medium of expression, such as on a server (17 U.S.C. § 102, a). Similar provisions exist in the intellectual property legislation of other countries, including those in the European Union (Commission, 2024). The owner of a copyright—which may be one or more individuals or corporate entities—has several exclusive rights under U.S. copyright law, including the rights to reproduce the work in copies, to create derivative works (works based on the copyrighted work), and to distribute the work, which also includes the right to authorize others to engage in such activities (17 U.S.C. § 106, [n. d.]). In addition, U.S. copyright law prohibits the “removal or alteration of copyright management information,” defined as including “the name of, and other identifying information about, the author of a work,” although the limits of this provision and its application to GenAI are as yet uncertain (17 U.S.C. § 1202, b).
The use of GenAI implicates these rights in several ways that are currently under consideration by the U.S. Copyright Office and the courts. (Office, 2023; The New York Times Company v. Microsoft Corporation, [n. d.]; Brittain, 2023; cop, 2023). For example, the plaintiffs in Doe v. GitHub (DOE 1 et al v. GitHub, [n. d.]), filed in November 2022, allege that Copilot removes copyright management information and violates the copyright licenses of the plaintiffs’ open source software by using that code as training data and generating code that is a near-identical reproduction of the plaintiffs’ code but without adhering to the terms of the associated licenses.
@@
The resolution of this and similar cases depends on both technological and legal questions. On the technological side, liability may depend on whether authors can successfully show that their works are indeed included in a particular LLM’s training data, as well as whether they can show that those works are reproduced or memorized in a way that would trigger liability under copyright law(Cooper and Grimmelmann, 2024). On the legal side, liability may depend on whether and how the generated content is similar to any particular author’s work. For example, generated content that is similar to previous works only with respect to the underlying concept or function is not infringing under U.S. copyright law because protection does not extend to any “idea, procedure, process, system, method of operation, concept, principle, or discovery” (17 U.S.C. §102, b). Additionally, some legal scholars argue that the use of a work protected by copyright as part of a model’s training data may be deemed a fair use under U.S. law and so not infringing (Lemley and Casey, 2020), similar to how Google’s use of copyrighted material in the database for its Google Books search engine was held to be fair use by the U.S. Court of Appeals for the Second Circuit in 2015 (Authors Guild, Inc. v. Google, Inc., 2015).
@@
Legislation both within and outside the U.S. also has implications for GenAI and copyright. Various U.S. states have attempted to enact legislation to govern GenAI (colorado-bill, 2024; california-regulation, 2024). In March 2024, the European Parliament approved the EU AI Act (Council of the European Union, 2024b), which, among other things, requires the providers of general-purpose AI models to “put in place a policy to comply with Union law on copyright and related rights, and in particular to identify and comply with, including through state-of-the-art technologies, a reservation of rights expressed pursuant to Article 4(3) of Directive (EU) 2019/790” (Council of the European Union, 2024a). This requires that creators of general-purpose AI models ensure upstream copyright compliance—apply due diligence in using copyrighted materials as part of their training data, obtain consent from copyright owners, and allow copyright holders to opt out—and ensure downstream copyright compliance, mitigating the risk—often due to overfitting—of generating verbatim content that is too similar to the (copyrighted) original. In 2023, an early announcement of the EU Global Principles for Artificial Intelligence, which led to the EU AI Act, encouraged some organizations, including French media organizations France Médias Monde and TF1, to forbid companies such as OpenAI from mining their data (Karathanasis, 2023); that year also saw a temporary ban on ChatGPT in Italy (McCallum, 2023). Even countries like Japan, which initially affirmed that training AI models on protected content was not a violation of copyright law (Lanz, [n. d.]), have started to consider additional restrictions designed to protect copyright holders (Shinbun, 2023). Gervais et al. (Gervais et al., 2024) provide an in-depth overview of the then-current copyright landscape around the globe.
2.2. Technical background
Additional work relevant to GenAI and copyright issues focuses on research that may help answer some of the technological questions raised above. For example, techniques exist to potentially reveal whether a machine learning model was trained on a certain data record, such as membership inference attacks (Duan et al., 2024; Hu et al., 2022; Majdinasab et al., 2024b; Zhang and Li, 2023; Z. Yang et al., 2023). Such attacks can extract training data from the target model’s output, including private and personal information (Carlini et al., 2021). Memorization of training data is a known issue, including in models trained on code (Yang et al., 2024), with research suggesting that typical measures aimed at preventing verbatim extraction of training data are not sufficient to prevent the data from being obtained with other methods, such as by utilizing style-transfer prompts (Ippolito et al., 2023). Tools like CopyrightCatcher (CopyrightCatcher, [n. d.]) have been developed to detect the presence of copyrighted material in model output, evidencing that this is an issue of concern to developers.
2.3. Related work
@@
Our work builds on this related work by contributing an in-depth exploration of developers’ perceptions of the copyright issues surrounding the use of AI-based code-generation tools, both in the U.S. and elsewhere around the world. Understanding these perspectives allows future regulation in this area to be informed by beliefs on the ground rather than in a vacuum. The U.S. Copyright Office, to “inform the Office’s study and help assess whether legislative or regulatory steps in this area are warranted,” has actively sought public comment on copyright issues pertaining to artificial intelligence (Office, 23). Even after receiving over 10,000 comments, the Copyright Office continues to study this developing area (Office, 2024b, 2025). In the same spirit, our work aims to capture the practices and views of developers to provide a resource for policymakers.
@@ 3. Study Methodology
The goal of this study is to investigate developers’ perceptions about licensing and copyright issues related to the use of GenAI technology for software development. (Although we use the term “software development” throughout the paper, our study primarily focuses on the use of GenAI for code generation as the main task that can lead to copyright and licensing issues.) The primary quality focus is to provide lawmakers, regulators, and other interested parties with a summary and analysis of relevant developer perspectives pertaining to the copyright and licensing questions that arise when developing software with the assistance of GenAI. The context consists of 574 developers of open-source projects hosted on GitHub, whom we surveyed via an online survey and follow-up interviews.
@@
    What other legal concerns do developers anticipate as the use of GenAI increases? There are a plethora of other potential legal issues to consider as GenAI becomes more widely adopted, including data privacy, generation of malicious content, and tort liability. In this RQ, we report developers’ perspectives on these topics which spread beyond copyright law.
@@ 3.1. Survey design
To design our survey, we followed general guidelines (Groves et al., 2009) as well as SE-specific best practices (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). The final questionnaire went through multiple iterations of review and improvements from six SE researchers and one law researcher, all authors of this paper. All researchers contributed to the design of the study, but in particular, the SE researchers contributed the domain experience to ask questions about procedures and practices relevant to developers, and the law researcher facilitated the identification of key areas that could be of concern related to copyright, licensing, etc., allowing the survey to ask the most pertinent questions possible. We carefully formulated the questions of the survey and ensured they were written clearly and concisely to avoid biasing respondents. We also conducted a pilot study with graduate students from our research lab who actively use GenAI tools. Based on their feedback, we further refined our questionnaire, improving questions that were ambiguous or hard to understand.
@@
  • The section on “Current use of GenAI for coding” (12 questions) asked about respondents’ experience using GenAI tools for code generation (C1, C2, C8), including the tools they used (C3); the development tasks for which they used the tools (C4); the benefits and challenges of using such tools (C5, C6); whether one examines the Terms of Service for tools (C7); organizational policies of using such tools (C9, C12); and procedures to document tool usage (C10, C11).
  • The section on “Understanding and Perception of Copyright Issues” (14 questions) asked about developers’ perceptions of copyright issues surrounding GenAI, including their views on if and how copyright, litigation, and legal risks factor into the use of GenAI (U1, U2, U3); the effects of these concerns on developers’ workflows (U4); whether the use of source code in training models should require attribution and/or monetary compensation (U6, U5); the use of their code in model training without permission (U7); having models produce code similar to theirs (U8); reusing without permission a prompt that they created (U9); ownership and copyrightability of generated code (U11, U12, U13); and other, broader concerns about copyright/legal issues stemming from development tasks with AI-based tools (U14).
  • The “Demographics” section (8 questions) asked about respondents’ age (D1); experience with SE and AI (D2, D3); their primary programming languages (D4); whether they develop open source or proprietary software (D5); their location (D6); and if they have received training on copyright (D6, D7).
The survey was designed to be completed in about 15 to 20 minutes and consisted mostly of multiple-choice questions (26), complemented by open-ended questions (7) that allowed participants to elaborate on their responses. Because we included logic in our survey to ask only those questions relevant to a participant’s indicated experience, not every participant saw every question. We included this logic to avoid asking questions about topics the participants indicated they did not have experience with. For example, if a participant indicated that they were not familiar with the practice of generating source code using GenAI (C1), we did not ask them questions about generating source code (C2-C12). The complete survey text and a brief description are found in our replication package (ano, 2024).